OpenStudy (anonymous):

multiply...(-3t)^2(2t^4)^2=-12t^4?

OpenStudy (slaaibak):

Close. 36t^10

OpenStudy (anonymous):

how did you get that?

OpenStudy (slaaibak):

Taking the first part: (-3t)^2 = (-3)^2 * (t)^2 = 9t^2 Taking the second part: (2t^4)^2 = (2)^2 * (t^4)^2 = 4t^8 Multiplying both parts: 9t^2 * 4t^8 = 9 * 4 * t^2 * t^8 = 36t^10